OpenStudy (anonymous):

2 divided by the 6th root of 27

OpenStudy (marigirl):

\[\frac{ 2 }{ \sqrt[6]{27} }\]

OpenStudy (marigirl):

\[\frac{ 2 }{6\sqrt{27}? }\]

OpenStudy (anonymous):

\[\frac{ 2 }{ \sqrt[6]{27} }\]\[\frac{2}{\sqrt[6]{3^{3}}}\]\[\frac{2}{\sqrt{3}}\]

OpenStudy (anonymous):

Or, another way: \[\frac{2}{\sqrt[6]{27}}\]\[\frac{2}{27^{\frac{1}6}}\]\[\frac{2}{(3^{3})^{\frac{1}6}}\]\[\frac{2}{3^{\frac{1}{2}}}\]\[\frac{2}{\sqrt{3}}\]

OpenStudy (anonymous):

Here's a picture if that's too small to read:

OpenStudy (anonymous):

yes i am new to this website so i was confused for a sec i was doing it off my phone

OpenStudy (anonymous):

mari girl it is the first one you sent me ! :)